Report of Activities on the RVIB N.B. Palmer
26 July 2001

Today (27 July) marks the official start of the second SO GLOBEC broad-scale survey. We arrived at station #1 in the early morning (~0600) and work commenced with a CTD, then a net tow, followed by deployment of BIOMAPER-II. We are currently headed between stations #3 and #4 and are position is -66° 00.226S; -69° 40.912W (at 1838, 27 July). Winds are out of the southwest (240) at 15 to 20 kts and the temperature is -12°C.

Yesterday (26 July) was our first full day in the pack ice and what a difference it made to the ride of the vessel. While there was some movement as the ship plowed through the ice flows, leaving a temporary trail of open water in its wake, there was no noticeable rolling or pitching. In the early afternoon, we stopped about 70 nm from Station #1 to conduct a test station where the instruments that will be routinely used during the grid survey were deployed into the water and samples for the laboratory analyses were obtained to enable the investigators to get their experimental procedures down and the kinks out. First in the water was the CTD, followed by the 1-m2 MOCNESS. After dinner, the ROV was deployed for under ice surveys of larval and adult krill. All tests were completed successfully, although the MOCNESS tow was done without the OPC or the Strobe Light being operational. Both instruments are still in a testing stage and were not ready for deployment on the net system. Once the gear was back on board, we continued steaming to Station #1. The trackline to the station was adjusted so that we steamed over SIO mooring #3 which was deployed last March on the SO GLOBEC mooring cruise. This passive listening equipment sitting on the seafloor requires periodic temperature profile measurements to be made so that sound velocity profiles can be computed. An XBT was done of the site as we passed over it and also a sonabuoy was deployed as described below.

John Klinck reports that the major activity of the CTD group was to continue with the XBT survey during the transit to the first station. We launched 22 XBT with 11 failures. The large failure rate is due partly to the use of old T5 probes and partly to the ice, which cuts the wire before the probe has a chance to get very deep. Nevertheless, we were able to complete the temperature survey with results that can be interpreted. We also got a temperature record near acoustic mooring SIO-3.

A re-analysis of the temperature records across the Drake Passage results in the following interpretation. We were in the vicinity of the Polar front since passing the 200-mile Argentine zone and crossed the front at a shallow angle. The traditional measure of the Polar front is the decent of the 2°C isotherm below 200 m. However, this isotherm has been waving from 150 to 300 m throughout our transit. Using a more holistic analysis of the whole record, it appears that we crossed the polar front for the first time around 60°S. Due to meandering of the front and our shallow crossing angle, we crossed again (northward) at 60.5°S and (southward) at 61.5°S. An additional meander between 62 and 63°S is indicated, but we only grazed the southern side of the front. As of the farthest south station, there is still no clear indication that we have crossed the Southern ACC front. The shallowing of the temperature maximum layer and the reduction in the temperature of temperature maximum to about 2°C indicates that we are close to the SACC front and are likely to have crossed it by the time we do the first CTD station.

We did a practice CTD station at 64° 39S; 69° 20W. The usual confusion and hesitation occurred, but the cast to 500 m was completed successfully. We drew water samples at 12 depths for analysis of oxygen, salinity, chlorophyll, and nutrients. We closed all bottles as a test of the system and to check for leaks.

As a minor comment regarding the met sensors, there was some interest in the air temperature which remained at -9.9°C throughout most of the day. Such steadiness in the atmosphere was not expected, especially since the wind chill continued to plunge through the day. A formatting error in the software code used to process the data was detected and corrected. Some of the recorded temperatures for today are incorrectly high by a degree or two.

Chris Ribic and Erik Chapman reported that on 26 July, as we traveled through densely packed first year floes interrupted by open water every kilometer or so, few birds were seen in the thick ice outside of two Adélie penguins that were resting on floes. In the open water, we saw a larger number of birds flying along the edges of ice and over grease ice forming in the leads. Here, mostly Snow and Antarctic Petrels were seen, but also a single Blue petrel and an immature Southern Giant Petrel were observed.

Ari Friedlaender noted that we have reached the pack ice, fully. After entering bits of brash last night, this morning found the ship well into 9-10/10ths small first year floes, pancakes, grey-white nilas, and brash. Large pond-shaped leads were scattered about in moderate frequency, but the concentrations of ice around them was constant. Marine mammal observations began at 1000 under overcast skies, no swell, and 25-30 knot winds out of the SW. The ship was making 7-8 knots through the ice en route to the first sampling station until stopping at 1250 for practice deployment of gear. Up to then, no sightings were made. Position at 1250: 64° 38.94S, 69° 20.07W. Other than seabirds, the only sightings made were of two Adélie penguins at 1208 (64° 33.65S, 69° 18.00W).

Ann Sirovic deployed 2 difar sonobouys on 26 July to listen for marine mammal calls. The first one was deployed a few miles before we stopped for instrument testing. It was monitored for almost 4 hours, but no biological sounds were heard. The second one was deployed as we were passing over SIO mooring #3 (-64° 59.406S; -69° 28.795W). This one had a lot of noise in the 20-40 Hz band, but still no biological sounds. We also dropped eight weighted light bulbs, starting over the sea floor instrument and then every 1 km as we steamed on towards Station #1. The light bulbs were intended to make a large broad-band noise when they imploded at depth and the noise recorded on the bottom moorings instrumentation and by the sonabuoy. None of the expected implosions were heard on the buoy, so this will have to be investigated in the next few days.

Carin Ashjian reports that the first 1-m2 MOCNESS tow down to 250 meters was conducted at the test station. The abundance of zooplankton collected was very low. At depth, there was very little biomass with only a few medium-sized euphausiids and a scant number of copepods from 200-250 meters. Very low abundance continued at shallower depths with some copepods, occasional ctenophores, and salps until 75 m. From 50-75 m, salps (10-20) and larval krill were more abundant. The 25-50 m interval was marked by 10-20 large ctenophores with purple combs. The top 25 m contained juvenile krill and a ctenophore and copepods, as well as plenty of ice.

Cheers, Peter